OpenStudy (anonymous):

find the distance between points a (-16,5) and b (-6, -5)

OpenStudy (saifoo.khan):

use the formula.

OpenStudy (anonymous):

I have the formula I am not sure once I get it to the radical of 200

OpenStudy (saifoo.khan):

u r correct. ^^

OpenStudy (anonymous):

simplify that radical, but yeah

OpenStudy (anonymous):

radical 10 and radical 20?

OpenStudy (saifoo.khan):

u will get, \[10\sqrt{2}\]

OpenStudy (saifoo.khan):

which is equal to ~14.14

OpenStudy (anonymous):

a - proj{a} b

OpenStudy (anonymous):

dot product over magnitudes

OpenStudy (anonymous):

\[\sqrt{200} = \sqrt{100 \cdot 2} = \sqrt{10^2 \cdot 2} = 10 \cdot \sqrt{2} = 10 \sqrt{2}\]